# Confluent Vandermonde matrix and related topics

### v2

@@ 6 Hermite polynomials
The set of polynomials $\mathcal{H}_{\Lambda}$ is called the Hermite polynomials of $\Lambda$. Given any polynomial $p(x)$ of degree less than $d$, one may recover the polynomial from the information $\operatorname{eval}_{\Lambda}(p(x))$ by
$[p(x)]=V_{\Lambda}^{-1}\operatorname{eval}_{\Lambda}(p(x))=H_{\Lambda}\operatorname{eval}_{\Lambda}(p(x))$
@@
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Suppose $p(x)\in{\mathbb{F}}[x]$ is a polynomial of degree less than $5$, and $p(1)=c_{1}$, $p(3)=c_{2}$, $\mathcal{D}^{(1)}(p(3))=c_{3}$, $p(6)=c_{4}$, $\mathcal{D}^{(1)}(p(6))=c_{5}$ are given. Then
@@
With $H_{\Lambda}=V_{\Lambda}^{-1}$, we have $H\operatorname{eval}_{\Lambda}(p(x))=[p(x)]$. Moreover, by Example 5.2, we may compute $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ and
$T^{-1}_{\Lambda}=\begin{bmatrix}\frac{1}{100}&0&0&0&0\\ 0&\frac{1}{18}&0&0&0\\ 0&\frac{1}{108}&\frac{1}{18}&0&0\\ 0&0&0&\frac{1}{45}&0\\ 0&0&0&-\frac{13}{675}&\frac{1}{45}\end{bmatrix}$
Let $\hat{h}_{1,0},\ldots,\hat{h}_{3,1}$ be as in Example 5.2. Since $H_{\Lambda}=V_{\Lambda}^{-1}=\hat{H}_{\Lambda}T^{-1}_{\Lambda}$, the Hermite polynomials are
$h_{1,0}(x)$ $=\frac{1}{100}\hat{h}_{1,0}(x)=\left[\frac{1}{100}\right](x-3)^{2}(x-6)^{2},$
$h_{2,0}(x)$ $=\frac{1}{18}\hat{h}_{2,0}(x)+\frac{1}{108}h_{2,1}(x)=(x-1)\left[\frac{1}{18}+\frac{1}{108}(x-3)\right](x-6)^{2},$
$h_{2,1}(x)$ $=\frac{1}{18}\hat{h}_{2,1}(x)=(x-1)\left[\frac{1}{18}(x-3)\right](x-6)^{2},$
$h_{3,0}(x)$ $=\frac{1}{45}\hat{h}_{3,0}(x)-\frac{13}{675}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left[\frac{1}{45}-\frac{13}{675}(x-6)\right],$
$h_{3,1}(x)$ $=\frac{1}{45}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left[\frac{1}{45}(x-6)\right].$
For each of these polynomials, its evaluations given by $\Lambda$ are all 00 except of a unique $1$. Therefore, for any given evaluations $c_{1},\ldots,c_{5}$, we have the Hermite interpolation
@@
Note that computing $T^{-1}_{\Lambda}$ is much easier than computing $V_{\Lambda}^{-1}$. In fact, the author of [13] used similar ideas to compute the inverse of a confluent Vandermonde matrix. The Hermite interpolation is also used to calculate the functions of companion matrices [5, 16].
@@ 7 Companion matrix
is the monic polynomial with roots in $\Lambda$, we write $C_{\Lambda}$ as the companion matrix of $p(x)$. Here, we always assume that the roots lie in the field. In particular, the discussion always work for an algebraically closed field.
@@
Brand[1] stated that
@@
is the Jordan canonical form of $C_{\Lambda}$ with exactly $1$ Jordan block for each eigenvalue. In particular, when all elements in $\Lambda$ are distinct, $V_{\Lambda}C_{\Lambda}V_{\lambda}^{-1}$ is a diagonal matrix.
@@
Let $f(x)=(x-\lambda_{1})(x-\lambda_{2})^{2}(x-\lambda_{3})^{2}$ such that $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Then $V_{\Lambda}C_{f}V_{\lambda}^{-1}$ will be a direct sum of $[\lambda_{1}],\begin{bmatrix}\lambda_{2}&0\cr 1&\lambda_{2}\end{bmatrix}$ and $\begin{bmatrix}\lambda_{3}&0\cr 1&\lambda_{3}\end{bmatrix}$.
This Jordan form result is elegant, but no proof was given in [1]. With the discussions in the previous sections, we can prove the Jordan form result above and gain better insight into it. Recall that $\operatorname{eval}_{\Lambda}(f(x))$ is a vector evaluating the polynomial $f(x)$. By definition, we have $\operatorname{eval}_{\Lambda}(p(x))=\mathbf{0}$. Moreover, each evaluation has the following properties.
@@
Let $f(x)\in{\mathbb{F}}[x]$ with distinct roots $\lambda_{1},\dots,\lambda_{q}$. Then
@@
$\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}(xf(x))\Big{|}_{x=\lambda}=\frac{1}{r!}\left[\binom{r}{1}\frac{d^{r-1}}{dx^{r-1}}(f(x))+x\cdot\frac{d^{r}}{dx^{r}}(xf(x))\right]\Big{|}_{x=\lambda}$
@@
$=\mathcal{D}^{(r-1)}(f(x))\big{|}_{x=\lambda}+\lambda\mathcal{D}^{(r)}(f(x))\big{|}_{x=\lambda}$
@@
Let $p_{j}(x)=x^{j}$. Then the columns of the $d\times d$ identity matrix $I$ are $[p_{0}(x)],\ldots,[p_{d-1}(x)]$ and
@@
Moreover, with our discussion in earlier sections, $V_{\Lambda}^{-1}$ comes from the Lagrange polynomials when all elements in $\Lambda^{-1}$ are distinct and from the Hermite polynomials in general. Let $\mathcal{E}_{d}=\{p_{0}(x),\ldots,p_{d-1}(x)\}$ and $\mathcal{H}_{\Lambda}$ the Hermite polynomials. Then both $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$ are bases of the space $\mathcal{P}_{d-1}$ of polynomials of degree less than $d$. Consider the linear transformation $\phi:\mathcal{P}_{d-1}\rightarrow\mathcal{P}_{d-1}$ defined by $\phi(f(x))=xf(x)\pmod{p(x)}$. Then $C_{\Lambda}$ can be viewed as the matrix representation of $\phi$ with respect to $\mathcal{E}_{d}$, while $J_{\Lambda}$ is the matrix representation of $\phi$ with respect to $\mathcal{H}_{\Lambda}$. Thus, $V_{\Lambda}$ and $H_{\Lambda}=V_{\Lambda}^{-1}$ are the change of basis matrices between $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$. This provides a comprehensive understanding of the Jordan decomposition of $C_{\Lambda}$.

### v3

@@ 6 Hermite interpolation polynomials
The set of polynomials $\mathcal{H}_{\Lambda}$ is called the Hermite interpolation polynomials of $\Lambda$. Given any polynomial $p(x)$ of degree at most d1d-1, one may recover the polynomial from the information $\operatorname{eval}_{\Lambda}(p(x))$ by
$p(x)]=V_{\Lambda}^{-1}\operatorname{eval}_{\Lambda}(p(x))=H_{\Lambda}\operatorname{eval}_{\Lambda}(p(x))$
@@
In general, computing $V_{\Lambda}^{-1}$ and formulating the Hermite interpolation polynomials are not straightforward; see, e.g., Respondek [24] for a survey. With Figure 1 in mind, finding the Hermite interpolation polynomials becomes easier with $V_{\Lambda}^{-1}=\hat{H}_{\Lambda}^{-1}T_{\Lambda}^{-1}$.
@@
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$ and $\mathbb{F}$ a field where $1,3,6$ are distinct. Suppose $p(x)\in{\mathbb{F}}[x]$ is a polynomial of degree at most $4$, and $p(1)=c_{1}$, $p(3)=c_{2}$, $\mathcal{D}^{(1)}(p(3))=c_{3}$, $p(6)=c_{4}$, $\mathcal{D}^{(1)}(p(6))=c_{5}$ are given. Then
@@
With $H_{\Lambda}=V_{\Lambda}^{-1}$, we have $H_{\Lambda}\operatorname{eval}_{\Lambda}(p(x))=[p(x)]$. Let $T_{\Lambda}^{-1}$ and $\hat{h}_{1,0},\ldots,\hat{h}_{3,1}$ be as in Example 5.3. Since $H_{\Lambda}=V_{\Lambda}^{-1}=\hat{H}_{\Lambda}T^{-1}_{\Lambda}$, the Hermite interpolation polynomials are
$h_{1,0}(x)$ $=\frac{1}{100}\hat{h}_{1,0}(x)=\left(\frac{1}{100}\right)(x-3)^{2}(x-6)^{2},$
$h_{2,0}(x)$ $=\frac{1}{18}\hat{h}_{2,0}(x)+\frac{1}{108}h_{2,1}(x)=(x-1)\left(\frac{1}{18}+\frac{1}{108}(x-3)\right)(x-6)^{2},$
$h_{2,1}(x)$ $=\frac{1}{18}\hat{h}_{2,1}(x)=(x-1)\left(\frac{1}{18}(x-3)\right)(x-6)^{2},$
$h_{3,0}(x)$ $=\frac{1}{45}\hat{h}_{3,0}(x)-\frac{13}{675}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left(\frac{1}{45}-\frac{13}{675}(x-6)\right),$
$h_{3,1}(x)$ $=\frac{1}{45}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left(\frac{1}{45}(x-6)\right).$
For each of these polynomials, its evaluations given by $\Lambda$ are all 0 except for a unique $1$. Therefore, for any given evaluations $c_{1},\ldots,c_{5}$, we have the Hermite interpolation
@@
Note that computing $T^{-1}_{\Lambda}$ is much easier than computing $V_{\Lambda}^{-1}$. In fact, the author of [15] used similar ideas to compute the inverse of a confluent Vandermonde matrix, and a faster algorithm was later proposed by Respondek [23]. The Hermite interpolation is also used to calculate the functions of companion matrices [7, 18].
@@ 7 Companion matrix
is the monic polynomial with roots in $\Lambda$, we write $C_{\Lambda}$ as the companion matrix of $p(x)$. Here, we always assume that the roots lie in the field. In particular, the discussion always works for an algebraically closed field.
@@
Aitken and Turnbull [2] stated that
@@
is the Jordan canonical form of $C_{\Lambda}$ with exactly $1$ Jordan block for each distinct eigenvalue. In particular, when all elements in $\Lambda$ are distinct, $V_{\Lambda}C_{\Lambda}V_{\lambda}^{-1}$ is a diagonal matrix.
@@
Let $\Lambda=\{\lambda_{1},\lambda_{2}^{(2)},\lambda_{3}^{(2)}\}$ and $p(x)=(x-\lambda_{1})(x-\lambda_{2})^{2}(x-\lambda_{3})^{2}$ such that $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Then $V_{\Lambda}C_{f}V_{\lambda}^{-1}$ will be a direct sum of $\lambda_{1}],\begin{bmatrix}\lambda_{2}&0\cr 1&\lambda_{2}\end{bmatrix}$ and $\begin{bmatrix}\lambda_{3}&0\cr 1&\lambda_{3}\end{bmatrix}$.
This Jordan form result is elegant, but no formal proof was given in [2]. With the discussions in the previous sections, we can prove the Jordan form result above and gain better insight into it. Recall that $\operatorname{eval}_{\Lambda}(f(x))$ is a vector evaluating the polynomial $f(x)$. By definition, we have $\operatorname{eval}_{\Lambda}(p(x))=\mathbf{0}$. Moreover, each evaluation has the following properties.
@@
Let $\Lambda$ be a multiset and $f(x)\in{\mathbb{F}}[x]$. Then
@@
$\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}(xf(x))\Big{|}_{x=\lambda}=\frac{1}{r!}\left(\binom{r}{1}\frac{d^{r-1}}{dx^{r-1}}(f(x))+x\cdot\frac{d^{r}}{dx^{r}}(xf(x))\right)\Big{|}_{x=\lambda}$
@@
$=\mathcal{D}^{(r-1)}(f(x))\big{|}_{x=\lambda}+\lambda\mathcal{D}^{(r)}(f(x))\big{|}_{x=\lambda}.$
@@
Let $p_{j}(x)=x^{j}$. Then the columns of the $d\times d$ identity matrix $I$ are $p_{0}(x)],\ldots,[p_{d-1}(x)]$ and
@@
Moreover, with our discussion in earlier sections, $V_{\Lambda}^{-1}$ comes from the Lagrange interpolation polynomials when all elements in $\Lambda^{-1}$ are distinct and from the Hermite interpolation polynomials in general. Let $\mathcal{E}_{d}=\{p_{0}(x),\ldots,p_{d-1}(x)\}$ and $\mathcal{H}_{\Lambda}$ the Hermite interpolation polynomials. Then both $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$ are bases of the space $\mathcal{P}_{d-1}$ of polynomials of degree at most d1d-1. Consider the linear transformation $\phi:\mathcal{P}_{d-1}\rightarrow\mathcal{P}_{d-1}$ defined by $\phi(f(x))=xf(x)\pmod{p(x)}$. Then $C_{\Lambda}$ can be viewed as the matrix representation of $\phi$ with respect to $\mathcal{E}_{d}$, while $J_{\Lambda}$ is the matrix representation of $\phi$ with respect to $\mathcal{H}_{\Lambda}$. Thus, $V_{\Lambda}$ and $H_{\Lambda}=V_{\Lambda}^{-1}$ are the change of basis matrices between $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$. This provides a comprehensive understanding of the Jordan decomposition of $C_{\Lambda}$.
@@
Acknowledgements
We thank the referee for their careful reading and helpful suggestions, which improved the quality of the paper. Chi-Kwong Li is an affiliate member of the Institute for Quantum Computing, University of Waterloo; his research was partially supported by the Simons Foundation Grant 851334. Jephian C.-H. Lin was supported by the National Science and Technology Council of Taiwan (grant no. NSTC-112-2628-M-110-003 and grant no. NSTC-113-2115-M-110-010-MY3).